# Omni-View: Unlocking How Generation Facilitates Understanding in Unified 3D Model based on Multiview images

### v1

@@ Abstract
This paper presents Omni-View, which extends the unified multimodal understanding and generation to 3D scenes based on multiview images, exploring the principle that “generation facilitates understanding”. Consisting of understanding model, texture module, and geometry module, Omni-View jointly models scene understanding, novel view synthesis, and geometry estimation, enabling synergistic interaction between 3D scene understanding and generation tasks. By design, it leverages the spatiotemporal modeling capabilities of its texture module responsible for appearance synthesis, alongside the explicit geometric constraints provided by its dedicated geometry module, thereby enriching the model’s holistic understanding of 3D scenes. Trained with a two-stage strategy, Omni-View achieves a state-of-the-art score of 55.4 on the VSI-Bench benchmark, outperforming existing specialized 3D understanding models, while simultaneously delivering strong performance in both novel view synthesis and 3D scene generation.
@@ 1 Introduction
Recently, unified multimodal models (UMMs) (team2024chameleon; xie2024show; deng2025bagel) have emerged as a pivotal area of research. The primary goal is to empower multimodal large language models (MLLMs) to both understand and generate visual signals present in our world, laying the foundation for artificial general intelligence.
Numerous methods (team2024chameleon; wu2024janus; wu2024vila; xie2024show; chen2025blip3; deng2025bagel) have achieved coexistence between 2D image understanding and generation. Among them, some studies (wu2024liquid; tong2024metamorph; yan2025can) aim to improve model performance by exploring the synergies arising from the interaction between understanding and generation. For example, Metamorph (tong2024metamorph) has extensively validated the role of understanding in improving generation performance. However, the potential and effectiveness of generation to improve understanding capabilities within unified models remain underexplored.
Due to the intrinsic geometric and spatiotemporal nature of 3D scenes and their multiview images, generative tasks such as geometry estimation and novel view synthesis are particularly well-suited for facilitating understanding in the 3D domain. This is attributable to the fact that 3D understanding tasks (yang2024thinking; zhang2025flatland) inherently necessitate robust geometric and spatiotemporal modeling capabilities, which can be effectively acquired through these generative tasks. Moreover, biomedical evidence (maus2013motion; nortmann2015primary) suggests that human understanding of 3D environments is governed by the capacity to generate and imagine future sensory and geometric data (keller2012sensorimotor; leinweber2017sensorimotor) in the observed scene. These findings provide us with a guidance: the “generation facilitates understanding” paradigm presents a promising approach to building a unified model for 3D scene understanding and generation.
@@
The architectural design aims at unifying 3D scene understanding and generation, leveraging geometry estimation and novel view synthesis to advance 3D scene understanding. Our Omni-View is built on Bagel (deng2025bagel), a strong unified framework in which the interaction between understanding and generation is facilitated by its shared multimodal self-attention. However, Bagel’s generative capacity is limited to RGB images, thus only capturing texture information, whereas 3D scene generation necessitates inclusion of both texture and geometric structure. In accordance with the dual generative objectives, the generation model in Omni-View is split into two distinct modules: a texture module and a geometry module. The texture module receives reference images, a list of targeted camera poses, and prompt tokens encoded by the understanding model to generate novel views of the scene. Meanwhile, the geometry module employs the hidden features from the understanding model and the latent output of the texture module to infer geometric information of novel views, such as depth maps and camera poses. This dual-pathway architecture empowers the model to develop both geometric and spatiotemporal modeling capabilities, which are essential for 3D scene understanding tasks.
@@
We evaluate Omni-View on scene understanding, spatial reasoning, and novel view synthesis tasks. The model achieves an impressive score of 55.4 on the VSI-Bench, exceeding current MLLMs designed for visual reasoning. It manifests particularly notable improvements in subtasks such as Relative Distance and Appearance Order, which require spatiotemporal modeling and the estimation of geometry acquired through generation tasks. It also exhibits superior performance compared to existing 3D understanding MLLMs in the area of 3D question answering (ma2023sqa3d; azuma2022scanqa). Furthermore, it effectively narrows the performance gap between unified models and specialized models focused on 3D visual localization (chen2020scanrefer). Furthermore, we attain robust results in the domain of novel view synthesis and scene generation, with particular emphasis on enhanced perceptual quality.
@@ 2 Related work
Scene understanding. Recent advances in understanding 3D scenes have been greatly provided by incorporating 3D or video input and 3D reconstruction prior. LLaVA-3D (zhu2024llava3d) and GPT4Scene (qi2025gpt4scene) function within the voxel space and BEV. Video3DLLM (zheng2024video3dllm) improves localization ability by encoding 3D coordinates as position embeddings, while Ross3D (wang2025ross3d) improves 3D understanding through visual-centric reconstruction. However, the dependency on 3D input of these methods poses practical application challenges. To alleviate this issue, VG-LLM (zheng2025learning) and SpatialMLLM (wu2025spatial) use the features of VGGT (wang2025vggt) as input, embedding the 3D piror in the model.
@@
Scene generation. A vital element of scene generation methods is the presence of an efficient proxy to represent the 3D scene, with panoramic images (team2025hunyuanworld), point clouds (yu2024viewcrafter), and Gaussian Splatting (yu2024wonderworld) among the viable options. In consideration of the advances in image and video diffusion models (rombach2022high; wan2025wan), contemporary 3D scene generation strategies reconstruct the scene’s geometry from a single view, employing conditioned video diffusion models to render the scene’s texture. ViewCrafter (yu2024viewcrafter) brings explicit 3D information (point cloud) to generation models through iterative reconstruction. WonderJourney (yu2024wonderjourney) generates a comprehensive set of view sequences. Voyager (huang2025voyager) reconstructs the scene from a single view and uses it as conditions for inpainting.
@@
Unified understanding and generation. In the domain of 3D scenes, a unified model for understanding and generation applicable in general scenarios remains absent. Hermes (zhou2025hermes) uses BEVs to design a unified understanding and generation model for autonomous driving. In contrast, significant progress has been made in 2D vision (zhang2025unified). These methods show primarily variations in architectures and training strategies. Chameleon (team2024chameleon) utilizes VQ-VAE for image tokenization, thereby improving generation competencies. However, its understanding capabilities are inferior to those of Janus (wu2024janus), which employs SigLIP (zhai2023sigmoid) as visual understanding encoder. VILA-U (wu2024vila) integrates understanding and generation within the image encoder, bypassing multitask gradient conflicts during MLLM training. BAGEL (deng2025bagel) implements task-based hard routing for MLLM, which also avoids gradient conflicts. Harmon (wu2025harmonizing) takes advantage of MAE’s reconstruction ability and downstream understanding enhancement. BLIP3o (chen2025blip3) introduces “understand first, then generation” training paradigm, thus achieving performance gains. Building on the progress in the 2D domain, this paper investigates a unified model in 3D scenes and explores how generation aids the understanding scheme within the framework.
@@ 3.1 Archtitectures
Figure 1: Architecture of Omni-View. Building upon Bagel (deng2025bagel), Omni-View consists of an understanding model and a generation model. The generation model is further composed of two specialized modules: one for texture and one for geometry. Trained via a two-stage process, Omni-View shows high effectiveness in scene understanding and novel view synthesis. Crucially, it unlocks the benefits of its generative capabilities to enhance the model’s understanding performance.
@@
Understanding Model. The understanding model is required to perform 3D scene or spatial comprehension tasks, including question answering, localization, and reasoning. This necessitates the model to extract semantic information from spatial-temporal data while measuring the geometric characteristics of the 3D scene. Specifically, given $n$ multiview images and a question $T_{ques}$, the understanding model produces the textual answer $T_{ans}$ through the next token prediction. In this process, the images are encoded by SigLIP (zhai2023sigmoid) and the text is tokenized by a frozen vocabulary $\tau(\cdot)$. Subsequently, tokens of multiview images and question $\tau(T_{ques})$ will be sequentially concatenated and fed into the understanding model. Moreover, the understanding model can also produce intermediate features, which will be used in the geometry module.
Texture Module. The texture module in the generation model is tasked with novel view synthesis using flow matching (lipman2022flow). This module processes a textual description $T_{des}$ and some reference images $I_{ref}$ of the current scene, together with a specified camera pose, to produce consistent novel views of this scene. Within this module, the reference images $I_{ref}$ are encoded using the FLUX-VAE encoder (flux2024) $\varepsilon(\cdot)$, and the vocabulary used to tokenize $T_{des}$ aligns with that used by the understandimg model. This process incorporates the camera pose control as delineated in MV-AR (hu2025auto), embedding the Plucker-Ray encoding $r_{i,j}=(o\times d,d)$ of the camera pose as the absolute position encoding in the model, where $o$ and $d$ donate the origin and direction of the ray, $(i,j)$ represents the pixel coordinate. This camera pose embedding exhibits adaptability to various image resolutions and demonstrates significant flexibility. The above process can be described as:
$F_{tex}=\text{Texture Module}\ ([\text{LM Head}(\tau(T_{des}));[\varepsilon(I_{ref});N_{tex}]+r]),$ (1)
@@
$F_{dep};\hat{g}]=\text{Geometry Module}\ ([F_{tex};N_{dep};q_{cam}],F_{und}),$ (2)
where $F_{dep}$ is the predicted depth noise of the geometry module. $\hat{g}$ is the predicted intrinsics and extrinsics of the camera, which is decoded by VGGT’s camera decoder (wang2025vggt).
@@ 3.2 Training Recipe
In contrast to the majority of novel view synthesis methods, the texture model employs a autoregressive generation framework. Specifically, during the generation of the $n$-th frame, the model is exposed to the visual data of the preceding n1n-1 frames, while excluding any subsequent frames. This autoregressive methodology enables the model to fully grasp the concept of temporal sequences, thereby enhancing its spatiotemporal modeling proficiency and effectively improving scene understanding. To optimize the 3D consistency of the sampled images, diffusion forcing (chen2024diffusion) is employed when training the texture module.
@@
In the texture module, the reference single-view image is used along with its depth map to reconstruct the initial point cloud of the scene. The images rendered from different views, projected through this point cloud, serve as conditions following Voyager (huang2025voyager). For the geometry module, it generates the depth map and camera pose from images synthesized by the texture module. Concurrently, it no longer relies on the features of the understanding model as conditions for cross-attention.
@@ 4.1 Experimental Setup
In our experiments, the understanding model and the texture module in the generation model are initialized using the pre-trained BAGEL-7B (deng2025bagel). The geometry module in the generation model is configured to have the same dimensions as the texture module, with a depth of four layers. For 3D scene understanding, a filtered dataset comprising 780k valid items, sourced from SQA3D (ma2023sqa3d), ScanQA (azuma2022scanqa), 3DOD (zheng2025learning), ScanRefer (chen2020scanrefer), VLM-3R (fan2025vlm), SPAR (zhang2025flatland) (234k subset) and llava-hound4 (zhang2024llavavideo) (64k subset), is meticulously curated for training. Regarding novel view synthesis, 61k video clips are carefully selected from re10k (zhou2018stereo). The corresponding depth maps are synthesized using the Voyager data pipeline (huang2025voyager) and captions are synthesized using the QwenVLMax (Bai2025Qwen25VLTR). During training, we do not use images from the scene understanding task to train the generation model to fully demonstrate that the understanding performance improvement brought by Omni-View does not come from memorizing the data in understanding tasks.
@@ 4.2 Main Comparisons
Benchmarks. We evaluate the 3D understanding performance of the model on question answering (ma2023sqa3d; azuma2022scanqa) and localization (chen2020scanrefer; zheng2025learning). During inference, we set the frame numbers as 32 following Video3DLLM (zheng2024video3dllm).
Comparison baselines. We compare Omni-View with models specifically designed for 2D or 3D visual understanding tasks (internvl2; wang2024qwen2vl; zhang2024llavavideo; huang2023leo; zhang2024chatscene; chen2024grounded; zhu2024llava3d; zheng2024video3dllm; qi2025gpt4scene; wang2025ross3d), as well as with some unified models applicable to video modalities (deng2025bagel). Unified models are evaluated after fine-tuning with the same data. Within the results table, we differentiate between models that incorporate explicit 3D input and those that do not. Although the incorporation of explicit 3D input improves model performance, it concurrently restricts applicability (zheng2024video3dllm).
@@
Table 1: Evaluation of 3D scene understanding. “–” indicates the number is not available for us. Bold and underline denote the best and second-best models without 3D scene input, respectively. For ScanRefer, the content in “()” indicates results without proposal refinement (zhang2025flatland).
Metrics. For SQA3D, we use the EM metric to evaluate the accuracy, which stands for top-1 exact match. EM-R means the refined EM following LEO (huang2023leo). For ScanQA, we use CIDEr (C), BLEU-4 (B-4), METEOR (M), ROUGE (R), and EM for more complete validation. For 3DOD, we use the average F1 score (F1) as a metric to assess the correspondence between the predicted and actual coordinates. For ScanRefer, we calculate the percentage of samples for which the Intersection over Union (IoU) exceeds thresholds of 0.25 and 0.5, respectively, between the predicted and true coordinates. The higher the above metrics, the better.
Results. As shown in Table 4.2.1, our analysis leads to four conclusions. (1) Our Omni-View exceeds all current MLLM methods that do not depend on 3D scene input. Within the SQA3D test set, Omni-View achieves an enhancement of 3.3 over SpatialMLLM in EM, while in the ScanQA validation set, it surpasses SpatialMLLM by an increment of 11.2 in CIDEr. For the object detection task, our unified model performs comparably with the 3D understanding model VG-LLM. (2) This performance improvement is mainly attributed to the architectural design and training scheme that we proposed. Compared with the fine-tuned BAGEL, our method still demonstrates substantial improvements. For example, on the SQA3D test set, Omni-View enhances EM by 2 points compared to the fine-tuned BAGEL (deng2025bagel); on the ScanQA validation set, Omni-View improves 7.5 in CIDEr. (3) The efficacy of our approach in the question-answering task is equivalent to advanced MLLM methods that require 3D scene input. In particular, in the ScanQA validation set, Omni-View achieved a performance comparable to Video3DLLM (zheng2024video3dllm) and LLaVA-3D (zhu2024llava3d). (4) However, there remains a notable disparity between methodologies that do not require 3D scene input and those that do, particularly in 3D grounding tasks. The experimental results of VG-LLM (zheng2025learning) also substantiate this observation.
@@
Benchmarks. We evaluate the 3D spatial reasoning ability of the model in VSI-Bench (yang2024thinking). During inference, we follow the VSI-Bench to set frame numbers ranging from 8 to 32 and frame resolution to 640p.
Comparison baselines. We compare our Omni-View with models specifically designed for 2D or 3D visual understanding tasks (Xue2024LongVILASL; Zhang2024LongCT; Bai2025Qwen25VLTR; ray2024sat; zhang2025flatland; zheng2025learning; wu2025spatial), as well as with some unified models applicable to video modalities (gpt4o; team2024gemini; deng2025bagel; xie2025show). Unified models are evaluated after fine-tuning with the same data we used.
@@
Results. The results on spatial reasoning tasks more fully demonstrate Omni-View’s improvement over previous methods in analyzing the relative or absolute position and orientation of spatial objects. With an average score of 55.4, our Omni-View ranks first among existing spatial reasoning MLLMs. Compared to existing Spatial-MLLM (wu2025spatial) and VG-LLM (zheng2025learning), Omni-View improves Spatial-MLLM by 11.6, 9.6, and 24.6 in Absolute Distance (Abs. Dist.), Room Size, and Relative Distance (Rel. Dist.), respectively. Omni-View improves VG-LLM by 9.8, 13.4, 25.1, 11.0, and 9.5 in Abs. Dist., Object Size (Obj. Size), Rel. Dist., Relative Direction (Rel. Dir.), and Appearance Order (Appr. Order), respectively. These tasks necessitate the model to thoroughly predict the spatiotemporal state and measure the geometric properties of the scene it observes (yang2024thinking). This demonstrates that our proposed method substantially enhances the model’s pertinent capabilities.
@@
Benchmarks. The model’s generation capacity is validated under two tasks: novel view synthesis (NVS) from a single view and scene generation. NVS from a single view necessitates that the model generate the subsequent 25 frames from the first image. 3D scene generation evaluates scenes that are reconstructed from the videos generated to 3DGS. The test scenes are randomly selected from the Re10k test set following chen2025flexworld.
Comparison baselines. We compare with scene generation methods (chung2023luciddreamer; wang2024motionctrl; ma2025you; yu2024viewcrafter; chen2025flexworld; huang2025voyager) and a unified model (deng2025bagel). To evaluate the performance of scene generation, we use Dust3R to reconstruct the generated videos, ensuring fairness in the comparison, according to chen2025flexworld. The evaluation is only conducted on the first 25 frames generated by the model following yu2024viewcrafter; zhai2025stargen.
NVS from single view and scene generation. Omni-View achieved the highest PSNR and SSIM, and lowest LPIPS score, indicating that its image quality could surpass that of other methods. However, in terms of pixel-level fidelity, Omni-View shows only slight improvements over popular scene generation models. This discrepancy may be attributed to Omni-View’s challenges in being precisely controlled via the camera pose. Visualization results are presented in the Appendix due to page limit.
@@
Table 3: Evaluation of novel view synthesis from single view and scene generation on Re10k.

### v2

@@ Abstract
This paper presents Omni-View, which extends the unified multimodal understanding and generation to 3D scenes based on multiview images, exploring the principle that “generation facilitates understanding”. Consisting of understanding model, texture module, and geometry module, Omni-View jointly models scene understanding, novel view synthesis, and geometry estimation, enabling synergistic interaction between 3D scene understanding and generation tasks. By design, it leverages the spatiotemporal modeling capabilities of its texture module responsible for appearance synthesis, alongside the explicit geometric constraints provided by its dedicated geometry module, thereby enriching the model’s holistic understanding of 3D scenes. Trained with a two-stage strategy, Omni-View achieves a state-of-the-art score of 55.4 on the VSI-Bench benchmark, outperforming existing specialized 3D understanding models, while simultaneously delivering strong performance in both novel view synthesis and 3D scene generation. The code and pretraiend models are open-sourced at https://github.com/AIDC-AI/Omni-View.
@@ 1 Introduction
Recently, unified multimodal models (UMMs) (Team, 2024a; Xie et al., 2024; Deng et al., 2025) have emerged as a pivotal area of research. The primary goal is to empower multimodal large language models (MLLMs) to both understand and generate visual signals present in our world, laying the foundation for artificial general intelligence.
Numerous methods (Team, 2024a; Wu et al., 2024a; c; Xie et al., 2024; Chen et al., 2025a; Deng et al., 2025) have achieved coexistence between 2D image understanding and generation. Among them, some studies (Wu et al., 2024b; Tong et al., 2024; Yan et al., 2025) aim to improve model performance by exploring the synergies arising from the interaction between understanding and generation. For example, Metamorph (Tong et al., 2024) has extensively validated the role of understanding in improving generation performance. However, the potential and effectiveness of generation to improve understanding capabilities within unified models remain underexplored.
Due to the intrinsic geometric and spatiotemporal nature of 3D scenes and their multiview images, generative tasks such as geometry estimation and novel view synthesis are particularly well-suited for facilitating understanding in the 3D domain. This is attributable to the fact that 3D understanding tasks (Yang et al., 2024; Zhang et al., 2025a) inherently necessitate robust geometric and spatiotemporal modeling capabilities, which can be effectively acquired through these generative tasks. Moreover, biomedical evidence (Maus et al., 2013; Nortmann et al., 2015) suggests that human understanding of 3D environments is governed by the capacity to generate and imagine future sensory and geometric data (Keller et al., 2012; Leinweber et al., 2017) in the observed scene. These findings provide us with a guidance: the “generation facilitates understanding” paradigm presents a promising approach to building a unified model for 3D scene understanding and generation.
@@
The architectural design aims at unifying 3D scene understanding and generation, leveraging geometry estimation and novel view synthesis to advance 3D scene understanding. Our Omni-View is built on Bagel (Deng et al., 2025), a strong unified framework in which the interaction between understanding and generation is facilitated by its shared multimodal self-attention. However, Bagel’s generative capacity is limited to RGB images, thus only capturing texture information, whereas 3D scene generation necessitates inclusion of both texture and geometric structure. In accordance with the dual generative objectives, the generation model in Omni-View is split into two distinct modules: a texture module and a geometry module. The texture module receives reference images, a list of targeted camera poses, and prompt tokens encoded by the understanding model to generate novel views of the scene. Meanwhile, the geometry module employs the hidden features from the understanding model and the latent output of the texture module to infer geometric information of novel views, such as depth maps and camera poses. This dual-pathway architecture empowers the model to develop both geometric and spatiotemporal modeling capabilities, which are essential for 3D scene understanding tasks.
@@
We evaluate Omni-View on scene understanding, spatial reasoning, and novel view synthesis tasks. The model achieves an impressive score of 55.4 on the VSI-Bench, exceeding current MLLMs designed for visual reasoning. It manifests particularly notable improvements in subtasks such as Relative Distance and Appearance Order, which require spatiotemporal modeling and the estimation of geometry acquired through generation tasks. It also exhibits superior performance compared to existing 3D understanding MLLMs in the area of 3D question answering (Ma et al., 2023; Azuma et al., 2022). Furthermore, it effectively narrows the performance gap between unified models and specialized models focused on 3D visual localization (Chen et al., 2020). Furthermore, we attain robust results in the domain of novel view synthesis and scene generation, with particular emphasis on enhanced perceptual quality.
@@ 2 Related work
Scene understanding. Recent advances in understanding 3D scenes have been greatly provided by incorporating 3D or video input and 3D reconstruction prior. LLaVA-3D (Zhu et al., 2024) and GPT4Scene (Qi et al., 2025) function within the voxel space and BEV. Video3DLLM (Zheng et al., 2024) improves localization ability by encoding 3D coordinates as position embeddings, while Ross3D (Wang et al., 2025a) improves 3D understanding through visual-centric reconstruction. However, the dependency on 3D input of these methods poses practical application challenges. To alleviate this issue, VG-LLM (Zheng et al., 2025) and SpatialMLLM (Wu et al., 2025a) use the features of VGGT (Wang et al., 2025b) as input, embedding the 3D piror in the model.
@@
Scene generation. A vital element of scene generation methods is the presence of an efficient proxy to represent the 3D scene, with panoramic images (Team et al., 2025), point clouds (Yu et al., 2024c), and Gaussian Splatting (Yu et al., 2024a) among the viable options. In consideration of the advances in image and video diffusion models (Rombach et al., 2022; Wan et al., 2025), contemporary 3D scene generation strategies reconstruct the scene’s geometry from a single view, employing conditioned video diffusion models to render the scene’s texture. ViewCrafter (Yu et al., 2024c) brings explicit 3D information (point cloud) to generation models through iterative reconstruction. WonderJourney (Yu et al., 2024b) generates a comprehensive set of view sequences. Voyager (Huang et al., 2025) reconstructs the scene from a single view and uses it as conditions for inpainting or restoration (Hu et al., 2025a; c; d).
@@
Unified understanding and generation. In the domain of 3D scenes, a unified model for understanding and generation applicable in general scenarios remains absent. Hermes (Zhou et al., 2025) uses BEVs to design a unified understanding and generation model for autonomous driving. In contrast, significant progress has been made in 2D vision (Zhang et al., 2025b). These methods show primarily variations in architectures and training strategies. Chameleon (Team, 2024a) utilizes VQ-VAE for image tokenization, thereby improving generation competencies. However, its understanding capabilities are inferior to those of Janus (Wu et al., 2024a), which employs SigLIP (Zhai et al., 2023) as visual understanding encoder. VILA-U (Wu et al., 2024c) integrates understanding and generation within the image encoder, bypassing multitask gradient conflicts during MLLM training. BAGEL (Deng et al., 2025) implements task-based hard routing for MLLM, which also avoids gradient conflicts. Harmon (Wu et al., 2025b) takes advantage of MAE’s reconstruction ability and downstream understanding enhancement. BLIP3o (Chen et al., 2025a) introduces “understand first, then generation” training paradigm, thus achieving performance gains. Building on the progress in the 2D domain, this paper investigates a unified model in 3D scenes and explores how generation aids the understanding scheme within the framework.
@@ 3.1 Archtitectures
Figure 1: Architecture of Omni-View. Building upon Bagel (Deng et al., 2025), Omni-View consists of an understanding model and a generation model. The generation model is further composed of two specialized modules: one for texture and one for geometry. Trained via a two-stage process, Omni-View shows high effectiveness in scene understanding and novel view synthesis. Crucially, it unlocks the benefits of its generative capabilities to enhance the model’s understanding performance.
@@
Understanding Model. The understanding model is required to perform 3D scene or spatial comprehension tasks, including question answering, localization, and reasoning. This necessitates the model to extract semantic information from spatial-temporal data while measuring the geometric characteristics of the 3D scene. Specifically, given $n$ multiview images and a question $T_{ques}$, the understanding model produces the textual answer $T_{ans}$ through the next token prediction. In this process, the images are encoded by SigLIP (Zhai et al., 2023) and the text is tokenized by a frozen vocabulary $\tau(\cdot)$. Subsequently, tokens of multiview images and question $\tau(T_{ques})$ will be sequentially concatenated and fed into the understanding model. Moreover, the understanding model can also produce intermediate features, which will be used in the geometry module.
Texture Module. The texture module in the generation model is tasked with novel view synthesis using flow matching (Lipman et al., 2022). This module processes a textual description $T_{des}$ and some reference images $I_{ref}$ of the current scene, together with a specified camera pose, to produce consistent novel views of this scene. Within this module, the reference images $I_{ref}$ are encoded using the FLUX-VAE encoder (Labs, 2024) $\varepsilon(\cdot)$, and the vocabulary used to tokenize $T_{des}$ aligns with that used by the understandimg model. This process incorporates the camera pose control as delineated in MV-AR (Hu et al., 2025b), embedding the Plucker-Ray encoding $r_{i,j}=(o\times d,d)$ of the camera pose as the absolute position encoding in the model, where $o$ and $d$ donate the origin and direction of the ray, $(i,j)$ represents the pixel coordinate. This camera pose embedding exhibits adaptability to various image resolutions and demonstrates significant flexibility. The above process can be described as:
$F_{tex}=TextureModule\ ([LM\text{-}Head(\tau(T_{des}));[\varepsilon(I_{ref});N_{tex}]+r]),$ (1)
@@
$F_{dep};\hat{g}]=GeometryModule\ ([F_{tex};N_{dep};q_{cam}],F_{und}),$ (2)
where $F_{dep}$ is the predicted depth noise of the geometry module. $\hat{g}$ is the predicted intrinsics and extrinsics of the camera, which is decoded by VGGT’s camera decoder (Wang et al., 2025b).
@@ 3.2 Training Recipe
In contrast to the majority of novel view synthesis methods, the texture model employs a autoregressive generation framework. Specifically, during the generation of the $n$-th frame, the model is exposed to the visual data of the preceding n1n-1 frames, while excluding any subsequent frames. This autoregressive methodology enables the model to fully grasp the concept of temporal sequences, thereby enhancing its spatiotemporal modeling proficiency and effectively improving scene understanding. To optimize the 3D consistency of the sampled images, diffusion forcing (Chen et al., 2024a) is employed when training the texture module.
@@
In the texture module, the reference single-view image is used along with its depth map to reconstruct the initial point cloud of the scene. The images rendered from different views, projected through this point cloud, serve as conditions following Voyager (Huang et al., 2025). For the geometry module, it generates the depth map and camera pose from images synthesized by the texture module. Concurrently, it no longer relies on the features of the understanding model as conditions for cross-attention.
@@ 4.1 Experimental Setup
In our experiments, the understanding model and the texture module in the generation model are initialized using the pre-trained BAGEL-7B (Deng et al., 2025). The geometry module in the generation model is configured to have the same dimensions as the texture module, with a depth of four layers. For 3D scene understanding, a filtered dataset comprising 780k valid items, sourced from SQA3D (Ma et al., 2023), ScanQA (Azuma et al., 2022), 3DOD (Zheng et al., 2025), ScanRefer (Chen et al., 2020), VLM-3R (Fan et al., 2025), SPAR (Zhang et al., 2025a) (234k subset) and llava-hound4 (Zhang et al., 2024c) (64k subset), is meticulously curated for training. Regarding novel view synthesis, 61k video clips are carefully selected from re10k (Zhou et al., 2018). The corresponding depth maps are synthesized using the Voyager data pipeline (Huang et al., 2025) and captions are synthesized using the QwenVLMax (Bai et al., 2025). During training, we do not use images from the scene understanding task to train the generation model to fully demonstrate that the understanding performance improvement brought by Omni-View does not come from memorizing the data in understanding tasks.
@@ 4.2 Main Comparisons
Benchmarks. We evaluate the 3D understanding performance of the model on question answering (Ma et al., 2023; Azuma et al., 2022) and localization (Chen et al., 2020; Zheng et al., 2025). During inference, we set the frame numbers as 32 following Video3DLLM (Zheng et al., 2024).
Comparison baselines. We compare Omni-View with models specifically designed for 2D or 3D visual understanding tasks (Team, 2024b; Wang et al., 2024a; Zhang et al., 2024c; Huang et al., 2023; Zhang et al., 2024a; Chen et al., 2024b; Zhu et al., 2024; Zheng et al., 2024; Qi et al., 2025; Wang et al., 2025a), as well as with some unified models applicable to video modalities (Deng et al., 2025). Unified models are evaluated after fine-tuning with the same data. Within the results table, we differentiate between models that incorporate explicit 3D input and those that do not. Although the incorporation of explicit 3D input improves model performance, it concurrently restricts applicability (Zheng et al., 2024).
@@
Table 1: Evaluation of 3D scene understanding. “–” indicates the number is not available for us. Bold and underline denote the best and second-best models without 3D scene input, respectively. For ScanRefer, the content in “()” indicates results without proposal refinement (Zhang et al., 2025a).
Metrics. For SQA3D, we use the EM metric to evaluate the accuracy, which stands for top-1 exact match. EM-R means the refined EM following LEO (Huang et al., 2023). For ScanQA, we use CIDEr (C), BLEU-4 (B-4), METEOR (M), ROUGE (R), and EM for more complete validation. For 3DOD, we use the average F1 score (F1) as a metric to assess the correspondence between the predicted and actual coordinates. For ScanRefer, we calculate the percentage of samples for which the Intersection over Union (IoU) exceeds thresholds of 0.25 and 0.5, respectively, between the predicted and true coordinates. The higher the above metrics, the better.
Results. As shown in Table 4.2.1, our analysis leads to four conclusions. (1) Our Omni-View exceeds all current MLLM methods that do not depend on 3D scene input. Within the SQA3D test set, Omni-View achieves an enhancement of 3.3 over SpatialMLLM in EM, while in the ScanQA validation set, it surpasses SpatialMLLM by an increment of 11.2 in CIDEr. For the object detection task, our unified model performs comparably with the 3D understanding model VG-LLM. (2) This performance improvement is mainly attributed to the architectural design and training scheme that we proposed. Compared with the fine-tuned BAGEL, our method still demonstrates substantial improvements. For example, on the SQA3D test set, Omni-View enhances EM by 2 points compared to the fine-tuned BAGEL (Deng et al., 2025); on the ScanQA validation set, Omni-View improves 7.5 in CIDEr. (3) The efficacy of our approach in the question-answering task is equivalent to advanced MLLM methods that require 3D scene input. In particular, in the ScanQA validation set, Omni-View achieved a performance comparable to Video3DLLM (Zheng et al., 2024) and LLaVA-3D (Zhu et al., 2024). (4) However, there remains a notable disparity between methodologies that do not require 3D scene input and those that do, particularly in 3D grounding tasks. The experimental results of VG-LLM (Zheng et al., 2025) also substantiate this observation.
@@
Benchmarks. We evaluate the 3D spatial reasoning ability of the model in VSI-Bench (Yang et al., 2024). During inference, we follow the VSI-Bench to set frame numbers ranging from 8 to 32 and frame resolution to 640p.
Comparison baselines. We compare our Omni-View with models specifically designed for 2D or 3D visual understanding tasks (Xue et al., 2024; Zhang et al., 2024b; Bai et al., 2025; Ray et al., 2024; Zhang et al., 2025a; Zheng et al., 2025; Wu et al., 2025a), as well as with some unified models applicable to video modalities (OpenAI, 2024; Team et al., 2024; Deng et al., 2025; Xie et al., 2025). Unified models are evaluated after fine-tuning with the same data we used.
@@
Results. The results on spatial reasoning tasks more fully demonstrate Omni-View’s improvement over previous methods in analyzing the relative or absolute position and orientation of spatial objects. With an average score of 55.4, our Omni-View ranks first among existing spatial reasoning MLLMs. Compared to existing Spatial-MLLM (Wu et al., 2025a) and VG-LLM (Zheng et al., 2025), Omni-View improves Spatial-MLLM by 11.6, 9.6, and 24.6 in Absolute Distance (Abs. Dist.), Room Size, and Relative Distance (Rel. Dist.), respectively. Omni-View improves VG-LLM by 9.8, 13.4, 25.1, 11.0, and 9.5 in Abs. Dist., Object Size (Obj. Size), Rel. Dist., Relative Direction (Rel. Dir.), and Appearance Order (Appr. Order), respectively. These tasks necessitate the model to thoroughly predict the spatiotemporal state and measure the geometric properties of the scene it observes (Yang et al., 2024). This demonstrates that our proposed method substantially enhances the model’s pertinent capabilities.
@@
Benchmarks. The model’s generation capacity is validated under two tasks: novel view synthesis (NVS) from a single view and scene generation. NVS from a single view necessitates that the model generate the subsequent 25 frames from the first image. 3D scene generation evaluates scenes that are reconstructed from the videos generated to 3DGS. The test scenes are randomly selected from the Re10k test set following Chen et al. (2025b).
Comparison baselines. We compare with scene generation methods (Chung et al., 2023; Wang et al., 2024b; Ma et al., 2025; Yu et al., 2024c; Chen et al., 2025b; Huang et al., 2025) and a unified model (Deng et al., 2025). To evaluate the performance of scene generation, we use Dust3R to reconstruct the generated videos, ensuring fairness in the comparison, according to Chen et al. (2025b). The evaluation is only conducted on the first 25 frames generated by the model following Yu et al. (2024c); Zhai et al. (2025).
NVS from single view and scene generation. Omni-View achieved the highest PSNR and SSIM, and lowest LPIPS score, indicating that its image quality could surpass that of other methods. However, in terms of pixel-level fidelity, Omni-View shows only slight improvements over popular scene generation models. This discrepancy may be attributed to Omni-View’s challenges in being precisely controlled via the camera pose. Visualization results are presented in the Appendix A.4.
@@
Table 3: Evaluation of novel view synthesis from single view and scene generation on Re10k. Bold and underline denote the best and second-best models, respectively.